# Resilience for Regular Path Queries: Towards a Complexity Classification

### v2

@@ 4.4. Hardness Gadgets
To show the hardness of resilience for other languages than $aa$, we will build hardness gadgets, or gadgets for brevity. Gadgets are used as building blocks to reduce from the vertex cover problem, generalizing the idea of Figure 3(a) from Section 4.1. We first define pre-gadgets as databases with two domain elements to which endpoint facts will be attached. We then define the encoding $\Xi$ of a directed graph $G$ by gluing together copies of the pre-gadgets following $G$, as in the proof of Proposition 4.1. We last define gadgets as pre-gadgets whose hypergraph of matches of the query $Q$ forms an odd chain. The vertex cover problem on $G$ then reduces to the resilience of $Q$ on $\Xi$.
@@
[…]
@@
[…]
We next explain how we use gadgets to encode directed graphs (see Figure 4 in Section 4.3):
@@
[…]
@@
{toappendix}
4.5. Proof of Section 4.5: Paths in Encodings
{claimrep}
[Paths in encodings] Let $G$ be a directed graph, let $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ be a pre-gadget, and let $\Xi$ be the encoding of $G$ with $\Gamma$. Then no walk of $\Xi$ passes through two internal elements belonging to different copies of $D$.
@@ 4.6. Condensed Hypergraph of Matches
Recall that a hitting set (also called hypergraph transversal (gottlob2004hypergraph)) of a hypergraph $\mathcal{H}=(V,E)$ is a subset $V^{\prime}\subseteq V$ such that $V^{\prime}\cap e\neq\emptyset$ for every $e\in E$. Thus, hitting sets are a generalization of vertex covers to hypergraphs. It is clear from the definition of resilience that $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)$ is equal to the minimum size of a hitting set of the hypergraph of matches $\mathcal{H}_{L,D}$. We now introduce two rewriting rules, called the condensation rules, that can be applied to any hypergraph $\mathcal{H}=(V,E)$ to simplify it without affecting the minimum size of hitting sets. For $v\in V$, write $E(v)=\{e\in E\mid v\in e\}$. The condensation rules are:
@@
If there are $v,v^{\prime}\in V$ such that $E(v)\subseteq E(v^{\prime})$, we can remove $v$ to obtain $\mathcal{H}^{\prime}=(V\setminus\{v\},\{e\setminus\{v\}\mid e\in E\})$. Intuitively, we can always replace $v$ by $v^{\prime}$ in hitting sets.
We say that $\mathcal{H}^{\prime}$ is a condensed hypergraph of $\mathcal{H}$ if $\mathcal{H}^{\prime}$ can be obtained from $\mathcal{H}$ by applying some sequence of condensation rules (note that we do not require that $\mathcal{H}^{\prime}$ cannot be further simplified using the rules). One can easily see that the rules preserve the minimum size of a hitting set:
{toappendix}
4.7. Proof of Section 4.7: Condensation Preserves Minimum Size of Hitting Sets
{claimrep}
[Condensation preserves minimum size of hitting sets] Let $\mathcal{H}^{\prime}$ be a condensed hypergraph of $\mathcal{H}$. Then the minimum size of a hitting set of $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are the same.
@@
By Section 4.7, we know that $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)$ is equal to the minimum size of a hitting set of any condensed hypergraph of $\mathcal{H}_{L,D}$, so we can always apply condensation rules in hardness proofs.
We can now define the requirement for a pre-gadget $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ to be a gadget [^5] Note that a notion generalizing gadgets has been introduced under the name of Independent Join Paths in (makhija2024unified), where it is used to show the hardness of resilience of conjunctive queries.: we require that the hypergraph of matches $\mathcal{H}_{L,D^{\prime}}$ of $L$ on the completion $D^{\prime}$ of $\Gamma$ can be condensed into an odd path. Formally, let $\mathcal{H}=(V,E)$ be a hypergraph and $u,v\in V$. We call $\mathcal{H}$ an odd path from $u$ to $v$ if all hyperedges of $\mathcal{H}$ have size 2 and the corresponding graph is an odd path from $u$ to $v$; i.e., there is a sequence $u=w_{1},\ldots,w_{2k}=v$ such that $E=\{\{w_{i},w_{i+1}\}\mid 1\leq i<2k\}$.
@@
The hardness pre-gadget whose completion is pictured in Figure 3(a) is a gadget for $aa$, as can be seen on the hypergraph of matches of Figure 3(b) (with no condensation rules applied).
@@
{toappendix}
4.8. Proof of Section 4.8: Gadgets Imply Hardness
{propositionrep}
[Gadgets imply hardness] Let $L$ be a reduced language. If there is a gadget $\Gamma$ for $L$, then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
{proofsketch}
The proof follows the same idea as that of Proposition 4.1. We reduce from the vertex cover problem: let $G^{\prime}$ be an undirected graph and $G$ be a directed graph obtained by choosing an arbitrary orientation for the edges of $G^{\prime}$. We build in linear time the encoding $\Xi$ of $G$ by $\Gamma$. We use the oracle to $\mathtt{RES}_{\mathrm{set}}(L)$ to obtain $r\coloneq\mathtt{RES}_{\mathrm{set}}(Q_{L},\Xi)$, which is equal to the minimum size of a hitting set of the hypergraph $\mathcal{H}_{L,\Xi}$ of matches of $L$ on $\Xi$. Using Section 4.5 and the fact that $\Gamma$ is a gadget, we show that there exists a condensation $\mathcal{H}^{\prime}$ of $\mathcal{H}_{L,\Xi}$ that has only hyperedges of size 2 and which (when seen as an undirected graph) is an odd-length subdivision $G^{\prime\prime}$ of $G^{\prime}$. By Section 4.7, $r=\mathtt{RES}_{\mathrm{set}}(Q_{L},\Xi)$ is equal to the minimum size of a hitting set of $\mathcal{H}^{\prime}$, i.e., $r$ is exactly the vertex cover number of $G^{\prime\prime}$, from which we recover the vertex cover number of $G^{\prime}$ via Section 4.2.
Hence, from now on, all hardness results will be shown by simply exhibiting hardness gadgets for the languages in question. We also wrote an implementation (gadgetrepo) to automatically check some correctness properties on our gadgets.
{toappendix}
To prove this, we reduce from the vertex cover problem. Let $G^{\prime}$ be an undirected graph, and let $G$ be an arbitrary orientation of $G$. We use the gadget to encode the input graph. The definitions that we impose on gadgets are designed to ensure that the encoding of $G$ has a condensed hypergraph of matches $\mathcal{H}^{\prime}$ which is an odd-length subdivision of the original graph $G^{\prime}$, so that by Section 4.2 computing a minimum vertex cover on $G^{\prime}$ reduces to computing a minimal hitting set of $\mathcal{H}^{\prime}$.
An odd subdivision of an undirected graph is an $\ell$-subdivision for some odd $\ell\in\mathbb{N}$. The technical claim used in the proof of Section 4.8 is the following:
{claim}
Let $G$ be a directed graph, let $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ be a gadget for $L$, and let $\Xi$ be the encoding of $G$ using $\Gamma$. Then $\mathcal{H}_{L,\Xi}$ has a condensed hypergraph that has only hyperedges of size $2$ and which is (when seen as an undirected graph) an odd subdivision of $G^{\prime}$, the undirected graph corresponding to $G$.
@@
Thanks to Section 4.5, one can check that the hypergraph of matches $\mathcal{H}_{L,\Xi}$ is precisely the union of the hypergraphs of matches of the various copies $D_{e}$ of $D$, but where in- and out-elements have been merged as in Definition 4.4. Then, considering the sequence of condensation rules to be done to obtain the condensed hypergraph of matches of the completion $D^{\prime}$ of $D$ that is an odd path as in Definition 4.6, we can do this sequence of simplifications on every copy $D_{e}$ of the gadget $D$ (noticing that by Definition 4.6 the source and target facts are not affected by the condensation rules), thus indeed obtaining a condensed hypergraph of matches of $\mathcal{H}_{L,\Xi}$ that is an odd subdivision of $G^{\prime}$. ∎
We can now prove Section 4.8 from Section 4.8:
Proof of Section 4.8.
We reduce from the vertex cover problem. Let $\Gamma$ be a gadget for $L$. Let $G^{\prime}$ be an undirected graph, and pick an arbitrary orientation $G$ of $G^{\prime}$. Construct the encoding $\Xi$ of $G$ by $\Gamma$: clearly this can be done in linear time. Use the oracle to compute the resilience of $Q_{L}$ on $\Xi$, written $r$, which is equal to the minimum size of a hitting set in $\mathcal{H}_{L,\Xi}$. By Section 4.8, $\mathcal{H}_{L,\Xi}$ has a condensed hypergraph $\mathcal{H}^{\prime}$ that is an odd subdivision of $G^{\prime}$ and moreover by Section 4.7 we have that the size of a minimum hitting set of $\mathcal{H}^{\prime}$ is also equal to $r$. We can then conclude using Section 4.2 (noticing that the length of the odd subdivision is a constant that depends only on $L$). ∎
@@
{toappendix}
4.10. Hypergraph of Matches for $axb|cxd$ (Proposition 4.8)
[…]
Figure 5. The hypergraph of matches of $axb|cxd$ on the completed gadget in Figure 3(c)
We use the gadget $\Gamma$ for the language $L=axb|cxd$ whose completion $D^{\prime}$ is depicted in Figure 3(c). Indeed, $\Gamma$ is clearly a pre-gadget. Further, we explain in Section 4.10 why the hypergraph of matches $\mathcal{H}_{L,D^{\prime}}$ of $L$ on $D^{\prime}$ (shown in Figure 5 in Section 4.10) can be condensed to the hypergraph of Figure 3(d), which is an odd path. Thus, $\Gamma$ is a gadget so Section 4.8 concludes. ∎
@@ 5. Hardness of Four-Legged Languages
In the previous section, we have presented hardness gadgets, and shown the hardness of resilience for the two specific languages $aa$ and $axb|cxd$. In this section we generalize the latter result to a subclass of the reduced non-local languages, called the four-legged languages: these languages intuitively feature a specific kind of counterexample to locality. We show that every four-legged language admits a gadget, so that resilience for such languages is always NP-hard by Section 4.8. We give two consequences of this result: it implies that resilience is hard for all non-star-free regular languages; and it also implies a dichotomy on resilience for languages featuring a neutral letter.
@@ 5.1. Four-Legged Languages
Recall the notion of local languages (Definition 3.1) from Section 3, which hinges on recognizability by certain automata (namely, local DFAs). We now give a known equivalent language-theoretic characterization of local languages, that we call being letter-Cartesian:
Definition 5.1 (Letter-Cartesian (sakarovitch2009elements)).
Note that, by exchanging the roles of $\alpha,\beta$ and $\gamma,\delta$, if $\alpha x\beta\in L$ and $\gamma x\delta\in L$ then also $\gamma x\beta\in L$. Hence the name letter-Cartesian: for each letter $x\in\Sigma$, the words of $L$ containing an $x$ can be expressed as a Cartesian product between what can appear left of an $x$ and what can appear right of an $x$. Formally, we have $L=\bigcup_{x\in\Sigma}L_{x}xR_{x}$ for some $L_{x},R_{x}\subseteq\Sigma^{*}$. It is known that letter-Cartesian languages are precisely the local languages: see (sakarovitch2009elements, Chp. II, Ex. 1.8, (d)) for a statement without proof, or Proposition 3.6 in Section 3.1 for a self-contained proof.
Example 5.2.
The local languages of Example 3.8 are letter-Cartesian. Take now $L=aa$. With $x\coloneq a$, $\alpha\coloneq a$, $\beta\coloneq\epsilon$, $\gamma\coloneq\epsilon$, $\delta\coloneq a$, we see that $L$ is not letter-Cartesian (as $aaa\notin L$), hence it is not local.
If a language $L$ is not letter-Cartesian, by definition there are $x\in\Sigma$ and $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ such that $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$. We then define the four-legged languages as those reduced languages that feature such a violation in which none of the words $\alpha$, $\beta$, $\gamma$, and $\delta$ is empty. Formally:
@@
A language $L$ is four-legged if it is reduced and has a letter $x\in\Sigma$ (the body) and four non-empty words $\alpha,\beta,\gamma,\delta\in\Sigma^{+}$ (the legs) with $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$.
@@
{toappendix}
5.2. Proof of Section 5.2: Hardness of Four-Legged Languages
{theoremrep}
[Hardness of four-legged languages] If $L$ is four-legged then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
{proofsketch}
Let $L$ be a four-legged language, with body $x\in\Sigma$ and legs $\alpha,\beta,\gamma,\delta\in\Sigma^{+}$ so that $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$. We first show that, up to changing the legs, we can ensure that no infix of $\alpha x\delta$ is in $L$. Then we distinguish two cases, depending on whether some infixes of $\gamma x\beta$ are in $L$ or not, and exhibit a gadget for each case. See Section 5.2 for details, with the gadgets and condensed hypergraphs of matches pictured in Figures 6 and 7.
{toappendix}
To show Section 5.2, we first show an intermediate lemma: for reduced languages that are four-legged, we can assume without loss of generality that the legs are stable, i.e., no infix of $\alpha x\delta$ is in $L$:
@@
{lemmarep}
Any reduced language which is four-legged with body $x\in\Sigma$ also has stable legs with body $x$.
@@
Let $L$ be a reduced language which is four-legged, and let $x\in\Sigma$ be a body and $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\Sigma^{+}$ be legs that witness it, i.e., $\alpha^{\prime}x\beta^{\prime}\in L$ and $\gamma^{\prime}x\delta^{\prime}\in L$ but $\eta^{\prime}\coloneq\alpha^{\prime}x\delta^{\prime}\notin L$.
If $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}$ are already stable (i.e., no infix of $\eta^{\prime}$ is in $L$) then we are done. Otherwise, as $\eta^{\prime}\notin L$, it means that there is a strict infix $\eta$ of $\eta^{\prime}$ that is in $L$. We know that $\eta$ must have a non-trivial intersection with $\alpha^{\prime}$, as otherwise $\eta\in L$ would be a strict infix of $\gamma^{\prime}x\delta^{\prime}\in L$, contradicting the fact that $L$ is reduced. Similarly, $\eta$ has a non-trivial intersection with $\delta^{\prime}$, as otherwise $\eta\in L$ would be a strict infix of $\alpha^{\prime}x\beta^{\prime}\in L$. So let us write $\alpha^{\prime}=\alpha_{2}\alpha_{1}$ and $\delta^{\prime}=\delta_{1}\delta_{2}$, with $\alpha_{1}$ and $\delta_{1}$ being non-empty, so that $\eta=\alpha_{1}x\delta_{1}$ is a strict infix of $\alpha^{\prime}x\delta^{\prime}$ that is in $L$. Note that $\alpha_{2}$ and $\delta_{2}$ cannot both be empty, as otherwise we have $\eta=\eta^{\prime}$ but we assumed that $\eta$ is a strict infix of $\eta^{\prime}$.

### v3

@@ 4.2. Hardness Gadgets
To show the hardness of resilience for other languages than $aa$, we will build hardness gadgets, or gadgets for brevity. Gadgets are used as building blocks to reduce from the vertex cover problem, generalizing the idea of Figure 3(b) from Section 4.1. We first define pre-gadgets as databases with two domain elements to which endpoint facts will be attached. We then define the encoding $\Xi$ of a directed graph $G$ by gluing together copies of the pre-gadgets following $G$, as in the proof of Proposition 4.1. We last define gadgets as pre-gadgets whose hypergraph of matches of the query $Q$ forms an odd path. The vertex cover problem on $G$ then reduces to the resilience of $Q$ on $\Xi$.
@@
A pre-gadget is a tuple $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ consisting of a database $D$, two distinguished domain elements $t_{\textup{in}}\neq t_{\textup{out}}$ of $D$ respectively called the in-element and out-element, and a label $a\in\Sigma$. We require that $t_{\textup{in}}$ and $t_{\textup{out}}$ never occur as heads of facts of $D$. The completion of $\Gamma$ is the database $D^{\prime}$ obtained from $D$ by adding two fresh distinct domain elements $s_{\textup{in}}$ and $s_{\textup{out}}$ and the two facts $s_{\textup{in}}\xrightarrow{a}t_{\textup{in}}$ and $s_{\textup{out}}\xrightarrow{a}t_{\textup{out}}$ using letter $a$.
@@
Figure 3(b) gives an example completion of a pre-gadget. The in-element and out-element of the pre-gadget are respectively $t_{u}$ and $t_{v}$, and the facts $s_{u}\xrightarrow{a}t_{u}$ and $s_{v}\xrightarrow{a}t_{v}$ are the facts added in the completion (so the pre-gadget has 4 edges and the completion has 6).
We next explain how we use gadgets to encode directed graphs (as in Figure 3(a)):
@@
    For each node $u$ in $G$, create a fact $s_{u}\xrightarrow{a}t_{u}$ in $\Xi$, where $s_{u}$ and $t_{u}$ are fresh domain elements.
@@
Claim 4.6 (Walks in encodings).
Let $G$ be a directed graph, let $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ be a pre-gadget, and let $\Xi$ be the encoding of $G$ with $\Gamma$. Then no walk of $\Xi$ passes through two internal elements belonging to different copies of $D$.
@@ 4.3. Condensed Hypergraph of Matches
Recall that a hitting set (also called hypergraph transversal (gottlob2004hypergraph, )) of a hypergraph $\mathcal{H}=(V,E)$ is a subset $V^{\prime}\subseteq V$ such that $V^{\prime}\cap e\neq\emptyset$ for every $e\in E$. Thus, hitting sets are a generalization of vertex covers to hypergraphs. It is clear from the definition of resilience that $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)$ is equal to the minimum size of a hitting set of the hypergraph of matches $\mathcal{H}_{L,D}$. We now introduce two rewriting rules, called the condensation rules, that can be applied to any hypergraph $\mathcal{H}=(V,E)$ to simplify it without affecting the minimum size of hitting sets. For $v\in V$, write $E(v)=\{e\in E\mid v\in e\}$. The condensation rules are:
@@
If there are $v,v^{\prime}\in V$ with $v\neq v^{\prime}$ such that $E(v)\subseteq E(v^{\prime})$, we can remove $v$ to obtain $\mathcal{H}^{\prime}=(V\setminus\{v\},\{e\setminus\{v\}\mid e\in E\})$. Intuitively, we can always replace $v$ by $v^{\prime}$ in hitting sets.
We say that $\mathcal{H}^{\prime}$ is a condensed hypergraph of $\mathcal{H}$ if $\mathcal{H}^{\prime}$ can be obtained from $\mathcal{H}$ by applying some sequence of condensation rules (note that we do not require that $\mathcal{H}^{\prime}$ cannot be further simplified using the rules). It can be remarked that the condensation rules are confluent in the sense that we always get isomorphic hypergraphs if we apply them to an input hypergraph until it is no longer possible, no matter the order in which we apply them: see (amarilli2025confluence, ). However, we will not need confluence for our purposes in the present paper. One can easily see that the rules preserve the minimum size of a hitting set:
Claim 4.8 (Condensation preserves minimum size of hitting sets).
Let $\mathcal{H}^{\prime}$ be a condensed hypergraph of $\mathcal{H}$. Then the minimum size of a hitting set of $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are the same.
@@
By 4.8, we know that $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)$ is equal to the minimum size of a hitting set of any condensed hypergraph of $\mathcal{H}_{L,D}$, so we can always apply condensation rules in hardness proofs.
We can now define the requirement for a pre-gadget $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ to be a gadget [^4] Note that a notion generalizing gadgets has been introduced under the name of Independent Join Paths in (makhija2024unified, ), where it is used to show the hardness of resilience of conjunctive queries.: we require that the hypergraph of matches $\mathcal{H}_{L,D^{\prime}}$ of $L$ on the completion $D^{\prime}$ of $\Gamma$ can be condensed into an odd path. Formally, let $\mathcal{H}=(V,E)$ be a hypergraph and $u,v\in V$. We call $\mathcal{H}$ an odd path from $u$ to $v$ if all hyperedges of $\mathcal{H}$ have size 2 and the corresponding graph is an odd path from $u$ to $v$; i.e., there is a sequence $u=w_{1},\ldots,w_{2k}=v$ such that $E=\{\{w_{i},w_{i+1}\}\mid 1\leq i<2k\}$.
@@
The hardness pre-gadget whose completion is pictured in Figure 3(b) is a gadget for $aa$, as can be seen on the hypergraph of matches of Figure 3(c) (with no condensation rules applied).
@@
Proposition 4.11 (Gadgets imply hardness).
Let $L$ be an infix-free language. If there is a gadget $\Gamma$ for $L$, then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
All our NP-hardness results in the paper will be proved using Proposition 4.11, by building gadgets for the queries that we consider. We note that we build the gadgets by hand, so it is possible that some queries may have shorter gadgets than the ones that we exhibit.
We prove Proposition 4.11 in the remaining of Section 4.3, by reducing from the vertex cover problem. Let $G^{\prime}$ be an undirected graph, and let $G$ be an arbitrary orientation of $G$. We use the gadget to encode the input graph. The idea is that the definitions that we impose on gadgets are designed to ensure that the encoding of $G$ has a condensed hypergraph of matches $\mathcal{H}^{\prime}$ which is an odd-length subdivision of the original graph $G^{\prime}$, so that by Proposition 4.2 and 4.8 computing a minimum vertex cover on $G^{\prime}$ reduces to computing a minimum hitting set of $\mathcal{H}^{\prime}$. Let us prove this formally.
We call an odd subdivision of an undirected graph an $\ell$-subdivision for some odd $\ell\in\mathbb{N}$. The technical claim used in the proof of Proposition 4.11 is the following:
Claim 4.12.
Let $G$ be a directed graph, let $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ be a gadget for $L$, and let $\Xi$ be the encoding of $G$ using $\Gamma$. Then there exists a condensed hypergraph of $\mathcal{H}_{L,\Xi}$ that has only hyperedges of size $2$ and which is (when seen as an undirected graph) an odd subdivision of $G^{\prime}$, the undirected graph corresponding to $G$.
@@
Thanks to 4.6, one can check that the hypergraph of matches $\mathcal{H}_{L,\Xi}$ is precisely the union of the hypergraphs of matches of the various copies $D_{e}$ of $D$, but where in- and out-elements have been merged as in Definition 4.5. Then, considering the sequence of condensation rules to be done to obtain the condensed hypergraph of matches of the completion $D^{\prime}$ of $D$ that is an odd path as in Definition 4.9, we can do this sequence of simplifications on every copy $D_{e}$ of the gadget $D$ (noticing that by Definition 4.9 the source and target facts are not affected by the condensation rules), thus indeed obtaining a condensed hypergraph of matches of $\mathcal{H}_{L,\Xi}$ that is an odd subdivision of $G^{\prime}$. ∎
We can now prove Proposition 4.11 from 4.12:
Proof of Proposition 4.11.
We reduce from the vertex cover problem. Let $\Gamma$ be a gadget for $L$. Let $G^{\prime}$ be an undirected graph, and pick an arbitrary orientation $G$ of $G^{\prime}$. Construct the encoding $\Xi$ of $G$ by $\Gamma$: clearly this can be done in linear time. Use the oracle to compute the resilience of $Q_{L}$ on $\Xi$, written $r$, which is equal to the minimum size of a hitting set in $\mathcal{H}_{L,\Xi}$. By 4.12, $\mathcal{H}_{L,\Xi}$ has a condensed hypergraph $\mathcal{H}^{\prime}$ that is an odd subdivision of $G^{\prime}$ and moreover by 4.8 we have that the size of a minimum hitting set of $\mathcal{H}^{\prime}$ is also equal to $r$. We can then conclude using Proposition 4.2 (noticing that the length of the odd subdivision is a constant that depends only on $L$). ∎
Hence, from now on, all hardness results in the paper will be shown by simply exhibiting hardness gadgets for the languages in question. We also wrote an implementation (gadgetrepo, ) to automatically perform some sanity checks on the gadgets that we construct. Namely, our implementation takes as input a description of a gadget along with a query. The query can be either explicit (e.g., $aa$, or $axb|cxd$), or it can be implicitly defined via requirements (e.g., words that are known to satisfy the query, or words for which it is known that no infix satisfies the query). The implementation then checks that the completion of the gadget satisfies the requirements outlined in Definition 4.3. Further, it exhaustively enumerates the directed paths in the gadget and checks which infixes of these paths are known to satisfy the query. This can be used to construct the condensed hypergraph of matches and check that it is indeed an odd path, to satisfy Definition 4.9. The implementation of (gadgetrepo, ) is also used to prepare the gadget drawings and condensed hypergraph of matches presented in the paper.
@@
By Proposition 4.11, it is enough to exhibit a gadget for $axb|cxd$. We claim that the pre-gadget $\Gamma$ whose completion $D^{\prime}$ is depicted in Figure 4(a) is a gadget. First, $\Gamma$ is clearly a pre-gadget, so we only need to show that it is a gadget, i.e., that it satisfies Definition 4.9.
We show in Figure 4(b) the hypergraph of matches $\mathcal{H}_{L,D^{\prime}}$ of the gadget whose completion $D^{\prime}$ is depicted in Figure 4(a). We now explain why $\mathcal{H}_{L,D^{\prime}}$ can be condensed to the hypergraph of Figure 4(c), which is an odd path. In this case, it suffices to apply node-domination rules to $\mathcal{H}_{L,D^{\prime}}$ to obtain $\mathcal{H}^{\prime}$ by eliminating all vertices corresponding to facts that are not endpoint facts and occur in only one hyperedge. For instance, we can first remove $t_{out}\xrightarrow{x}15$ because it is dominated by $15\xrightarrow{b}16$, then remove $12\xrightarrow{x}15$ because it is dominated by $13\xrightarrow{x}12$, and so on, until we obtain $\mathcal{H}^{\prime}$.
Thus, $\Gamma$ is indeed a gadget, which concludes the proof. ∎
@@ 5. Hardness of Four-Legged Languages
In the previous section, we have presented hardness gadgets, and shown the hardness of resilience for the two specific languages $aa$ and $axb|cxd$. In this section we generalize the latter result to a subclass of the infix-free non-local languages, called the four-legged languages: these languages intuitively feature a specific kind of counterexample to locality. We show that every four-legged language admits a gadget, so that resilience for such languages is always NP-hard by Proposition 4.11. We give two consequences of this result: it implies that resilience is hard for all non-star-free regular languages; and it also implies a dichotomy on resilience for languages featuring a neutral letter.
@@ 5.1. Four-Legged Languages
Recall the notion of local languages (Definition 3.1) from Section 3, which hinges on recognizability by certain automata (namely, local DFAs), and which was equivalent to the notion of being letter-Cartesian (Proposition 3.5).
If a language $L$ is not letter-Cartesian, by definition there are $x\in\Sigma$ and $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ such that $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$. We then define the four-legged languages as those infix-free languages that feature such a violation in which none of the words $\alpha$, $\beta$, $\gamma$, and $\delta$ is empty. Formally:
@@
A language $L$ is four-legged if it is infix-free and has a letter $x\in\Sigma$ (the body) and four non-empty words $\alpha,\beta,\gamma,\delta\in\Sigma^{+}$ (the legs) with $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$.
@@
Theorem 5.3 (Hardness of four-legged languages).
If $L$ is four-legged then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
We prove Theorem 5.3 in the remaining of Section 5.1. To this end, we first show an intermediate lemma: for languages that are four-legged, we can assume without loss of generality that the legs are stable, i.e., no infix of $\alpha x\delta$ is in $L$:
@@
Lemma 5.5.
Any infix-free language which is four-legged with body $x\in\Sigma$ also has stable legs with body $x$.
@@
Let $L$ be an infix-free language which is four-legged, and let $x\in\Sigma$ be a body and $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\Sigma^{+}$ be legs that witness it, i.e., $\alpha^{\prime}x\beta^{\prime}\in L$ and $\gamma^{\prime}x\delta^{\prime}\in L$ but $\eta^{\prime}\coloneq\alpha^{\prime}x\delta^{\prime}\notin L$.
If $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}$ are already stable (i.e., no infix of $\eta^{\prime}$ is in $L$) then we are done. Otherwise, as $\eta^{\prime}\notin L$, it means that there is a strict infix $\eta$ of $\eta^{\prime}$ that is in $L$. We know that $\eta$ must have a non-trivial intersection with $\alpha^{\prime}$, as otherwise $\eta\in L$ would be a strict infix of $\gamma^{\prime}x\delta^{\prime}\in L$, contradicting the fact that $L$ is infix-free. Similarly, $\eta$ has a non-trivial intersection with $\delta^{\prime}$, as otherwise $\eta\in L$ would be a strict infix of $\alpha^{\prime}x\beta^{\prime}\in L$. So let us write $\alpha^{\prime}=\alpha_{2}\alpha_{1}$ and $\delta^{\prime}=\delta_{1}\delta_{2}$, with $\alpha_{1}$ and $\delta_{1}$ being non-empty, so that $\eta=\alpha_{1}x\delta_{1}$ is a strict infix of $\alpha^{\prime}x\delta^{\prime}$ that is in $L$. Note that $\alpha_{2}$ and $\delta_{2}$ cannot both be empty, as otherwise we have $\eta=\eta^{\prime}$ but we assumed that $\eta$ is a strict infix of $\eta^{\prime}$.